OpenStudy (bilalkhaliq22):

evaluate 10 sigma n=2 -2n-3? how do you do this

OpenStudy (astrophysics):

Do the same thing I showed you earlier \[\sum_{n=2}^{10} -2n-3 \implies -2(2)-3+-2(3)-3+...\] start at n = 2 and end at n = 10

OpenStudy (bilalkhaliq22):

what is the + is for after 3?

OpenStudy (astrophysics):

That means the pattern continues, so do it up till 10

OpenStudy (astrophysics):

All you're doing is plugging in the "n's" into the equation

OpenStudy (astrophysics):

and sigma implies to sum it up, so you plug in n = 2 and keep plugging in n's up till 10, and then you add them all up

OpenStudy (bilalkhaliq22):

Is -135 the correct answer. that what I got when I add them all up

OpenStudy (astrophysics):

Good job!

OpenStudy (bilalkhaliq22):

evaluate 5 \[\sum_{n=1}^{5} 3(-2)^{n-1}\]

OpenStudy (astrophysics):

Same thing as before except now n is a exponent

OpenStudy (astrophysics):

Remember an exponent tells you how many times you multiply by itself and anything to the power of 0 = 1, \[x^0=1\]

OpenStudy (bilalkhaliq22):

so is it 3(-2)^1-1 3(-2)^2-1

OpenStudy (astrophysics):

yes exactly up to 5

OpenStudy (bilalkhaliq22):

I feel like I did something wrong because i add them up it is equal to -71

OpenStudy (bilalkhaliq22):

my options are a.−93 b.−33 c.33 d.93

OpenStudy (astrophysics):

\[3(-2)^{1-1} = 3(-2)^0 = 3(1) = 3\]

OpenStudy (astrophysics):

\[3(-2)^{2-1} = 3(-2)^1 = -6\]

OpenStudy (astrophysics):

\[3(-2)^{3-1} =3(-2)^2 = 3(4)=12\]

OpenStudy (astrophysics):

now do n = 4 and n = 5

OpenStudy (bilalkhaliq22):

\[3(-2)^{4-1} = 3(-2)^{3} =?\]

OpenStudy (bilalkhaliq22):

@Astrophysics

OpenStudy (bilalkhaliq22):

@Ashleyisakitty

OpenStudy (bilalkhaliq22):

@Agl202

OpenStudy (bilalkhaliq22):

got it thank you very much

OpenStudy (bilalkhaliq22):

\[\sum_{n=3}^{12} 20(0.5)^{n-1}\]